Imagine a space elevator with m = 30 kg/m extending from the surface (6400 km) to the geosynchronous height of 40,000 km. Take g to be changing with the height. If a stray satellite hits the top of the elevator find the time that the wave will take to reach the surface. The answer should be a few hours or so.
1
Expert's answer
2020-12-21T11:35:05-0500

"m = 30kg/m\\
R = 6400 km\\
h = 40,000km\\"


"\begin{aligned}
g_n &= g(1+\frac{h}{R_e})^{-2}\\
&= 9.8(1+\frac{40000}{6400})^{-2}\\
&= 0.186
\end{aligned}"


"\begin{aligned}
T &= \dfrac{2π}{R_e}\sqrt{\dfrac{(R_e+h)^3)}{g}}\\
&= \dfrac{2π}{6400×10^3}\sqrt{\dfrac{((40000+6400)×10^3)^3)}{g}}\\
&= 99 × 10^3 sec
\end{aligned}"


"\begin{aligned} 
v &= \dfrac{2πr}{T}\\ \\
&= \dfrac{2π(R_e + h)}{T}\\ \\
&= \dfrac{2π((40000+ 6400)×10^3)}{99×10^3}\\ \\
&= 2.9×10^3ms^{-1}
\end{aligned}"


"\begin{aligned}
t &= \dfrac{s}{v_0}\\ \\
&= \frac{40000×10^3}{2.9×10^3} = 13,793 sec
\end{aligned}"


converting 13793 seconds to hour, we divide the number of seconds by 3600


"\dfrac{13793}{3600} = 3.83 \ hours ="


"3\ hours \ \ 50\ minutes"


Therefore, it will take the wave 3 hours 50 minutes to reach the surface.